OpenStudy (shootingstar28):

The graph of g(x) is the graph of f(X) = x + 6 reflected across the x-axis. Which equation describes g? Question 3 options: g(x) = -x + 6 g(x) = -6x - 6 g(x) = -x - 6 g(x) = x - 6

OpenStudy (3mar):

May I help?

OpenStudy (shootingstar28):

Go ahead :'3

OpenStudy (3mar):

Thank you.

OpenStudy (3mar):

Reflection across x-axis: (x, y) > (x, -y)

OpenStudy (3mar):

@ShootingStar28

OpenStudy (tgstudios):

from that formula I would think D, because the y-intercept would become a negative.

OpenStudy (3mar):

Not a direct answer!

OpenStudy (tgstudios):

ummm.. it wasn't to direct, but sorry.. :/

OpenStudy (tgstudios):

that is in the code of conduct... sry @3mar

OpenStudy (tgstudios):

but do you understand @ShootingStar28 ?

OpenStudy (shootingstar28):

No.

OpenStudy (3mar):

Where are you stuck, @ShootingStar28?

OpenStudy (shootingstar28):

Um the whole thing...

OpenStudy (3mar):

Step by step

OpenStudy (3mar):

This is f(x) with no reflection!

OpenStudy (shootingstar28):

I guess like D?

OpenStudy (3mar):

I need you get the idea first then choose the answer

OpenStudy (3mar):

Apply the rule of Reflection across x-axis:(x, y) > (x, -y) but we have y=x+6 and : y would be -y so that -y=x+6 which means y=-(x+6) y=-x-6

OpenStudy (shootingstar28):

So the answer is C instead?

OpenStudy (3mar):

Yes, but Did you get it?

OpenStudy (shootingstar28):

Yeah

OpenStudy (3mar):

Happy to hear that!

OpenStudy (3mar):

Any more questions?